### v2

# Christopher P. Chambers and Christopher Turansick

Abstract
We study identification and linear independence in random utility models. We characterize the dimension of the random utility model as the cyclomatic complexity of a specific graphical representation of stochastic choice data. We show that, as the number of alternatives grows, any linearly independent set of preferences is a vanishingly small subset of the set of all preferences. We introduce a new condition on sets of preferences which is sufficient for linear independence. We demonstrate by example that the condition is not necessary, but is strictly weaker than other existing sufficient conditions.
1 Introduction
The random utility model is the basis for much of modern discrete choice analysis. Random utility moves beyond the classic rational model and supposes that there is some distribution over preferences inducing choice. This distribution over preferences is typically interpreted as heterogeneity within a population or variation of a single agent’s preference across time. A common desideratum for a model in empirical practice is identification as identification allows for proper counterfactual analysis. The random utility model is known to be unidentified, however there are many refinements of the random utility model which recover identification. We study those models who recover identification by limiting the set of allowable preferences in the model.
Identifying assumptions by nature put restrictions on a model in order to recover identification. In the case of models which assume away certain preferences, this is equivalent to assuming away certain types of behavior. Our first question is how much behavior is or how many preferences are assumed away when we recover identification. Our main result shows that the maximum number of preferences in a linearly independent set of preferences is equal to $(n-2)2^{n-1}+2$, where $n$ is the number of alternatives in our environment. It follows that the ratio of preferences in a model identified through preference restrictions and total number of preferences given a set of alternatives of size $n$ scales at a rate of $\mathcal{O}(\frac{2^{n-1}}{(n-1)!})$. This tells us that, as $n$ grows large, the number of preferences in a model identified through preference restrictions is vanishingly small when compared to the total number of possible preferences. When $n=5$, already more than half of the preferences become unavailable in an identified model. When $n=9$, more than $99.5\%$ of the preferences must be discarded in order to write an identified model. Thus identifying assumptions of this nature must assume away almost all types of behavior.
Our second goal in this paper is to offer a new identifying restriction for random utility models. Our new identifying restriction is called edge decomposability. In our setup a random utility model is simply a collection of feasible preferences. Edge decomposability asks that, for a given model, for every non-empty subset of preferences of that model there exists some preference in that subset with a contour set unique to that preference within that subset. More specifically it asks that for every non-empty subset of preferences, there exists some pair $(x,A)$ where $x$ is an alternative in menu $A$ such that there is a single preference in this subset that chooses $x$ from $A$ but fails to choose $x$ from any strict superset of $A$. While edge decomposability is a sufficient condition for linear independence of our set of preferences, surprisingly it fails to be a necessary condition. We then compare edge decomposability with the identifying restrictions of Apesteguia et al. (2017) and Turansick (2022). We show that every set of preferences which can be the support of some distribution over preferences satisfying the conditions of either Apesteguia et al. (2017) or Turansick (2022) must also be edge decomposable. Further, we show that there are sets of preferences that are edge decomposable yet cannot be the support of a distribution satisfying either the conditions of Apesteguia et al. (2017) or Turansick (2022).
The rest of this paper proceeds as follows. In Section 2 we introduce our setup and some preliminary constructions. In Section 3 we provide our main result and discuss its proof. In Section 4 we introduce our identifying restriction and compare it to the restrictions of Apesteguia et al. (2017) and Turansick (2022). Section 5 provides a complete classification of all maximal identified models for the case of $n=4$ alternatives. Finally, we conclude with a discussion of the related literature in Section 6.
2 Model and Preliminaries
We work with a finite set of alternatives $X_{n}=\{1,\dots,n\}$ where each alternative is indexed by a number between $1$ and $n$. Further, we use $x$ and $y$ to denote arbitrary elements of $X_{n}$. We are interested in strict preferences over this environment. Denote the set of linear orders over $X_{n}$ by $\mathcal{L}(X_{n})$. [^2] A linear order is complete, antisymmetric, and transitive. A typical linear order is denoted by $\succ$.
Definition 2.1.
A model on $X_{n}$ (or simply a model) is a nonempty set of linear orders $M$: $\emptyset\subsetneq M\subseteq\mathcal{L}(X_{n})$.
We denote the collection of all models on $X_{n}$ as $\mathcal{M}_{n}$. A model restricts which preferences an agent may have. Using the standard $\Delta$ notation to represent the set of probability measures over a set, let $\nu\in\Delta(M)$ denote a typical probability distribution over the preferences in model $M$. Further, let $2^{X}$ denote the power set of $X$ and let $A$ denote a typical element of $2^{X}$.
Definition 2.2.
A function $p:X_{n}\times 2^{X_{n}}\setminus\{\emptyset\}\rightarrow\mathbb{R}$ is a random choice rule if it satisfies the following.
  • $p(x,A)\geq 0$ for all $x\in A$
  • $\sum_{x\in A}p(x,A)=1$
Random choice rules correspond to our data primitive. They denote how frequently an alternative $x$ is chosen from a menu $A$. Given $\nu\in\Delta(\mathcal{L}(X_{n}))$, let $p_{\nu}$ denote the random choice rule induced by $\nu$. Specifically,
$p_{\nu}(x,A)=\sum_{\succ\in\mathcal{L}(X_{n})}\nu(\succ)\mathbf{1}\{x\succ y\text{ }\forall y\in A\setminus\{x\}\}.$ (1)
Definition 2.3.
A model $M$ is identified if for any $\nu,\nu^{\prime}\in\Delta(M)$, $p_{\nu}=p_{\nu^{\prime}}$ implies that $\nu=\nu^{\prime}$.
Given $X_{n}$, we denote the set of identified models as $\mathcal{I}_{n}$. Our goal is to provide a tight upper bound on the size of models in $\mathcal{I}_{n}$. Specifically, our main question is what is the largest number of preferences that admit an identified model. In order to answer this question, we rely on the Möbius inverse of a random choice rule. [^3] The Möbius inverse is a fairly general method of inverting cumulative sums on certain partially ordered sets. In our case, the partially ordered set under consideration is $(X_{n},\subseteq)$, and the inversion formula is referred to as the inclusion-exclusion principle. A classical reference is Rota (1964).
Definition 2.4.
The Möbius inverse of $p(x,A)$ is a function $q:X\times 2^{X}\setminus\{\emptyset\}\rightarrow\mathbb{R}$ defined as follows.
$\begin{split}q(x,A)&=p(x,A)-\sum_{A\subsetneq B}q(x,B)\\ &=\sum_{A\subseteq B}(-1)^{|B\setminus A|}p(x,B)\end{split}$ (2)
The Möbius inverse $q(x,A)$ simply keeps track of how much choice probability is being added to or removed from the choice of $x$ at menu $A$. Further, the Möbius inverse function has a special relationship with random choice rules induced by models.
Definition 2.5.
A random choice rule $p$ is stochastically rational if there exists some $\nu\in\Delta(\mathcal{L}(X))$ such that $p=p_{\nu}$.
Consider the set of preferences which choose $x$ from $A$ but do not choose $x$ from any superset of $A$. This set is given by $L(x,A)=\{\succ|z\succ x\succ y\text{ }\forall y\in A\setminus\{x\},\text{ }\forall z\in X\setminus A\}$.
Theorem 2.1 (Falmagne (1978)).
Let $q$ be the Möbius inverse of $p_{\nu}$. Then for all $A\subseteq X$ and all $x\in A$, $q(x,A)=\nu(L(x,A))$.
This result tells us that $q(x,A)$ is the probability under $p_{\nu}$ of $L(x,A)$. Further, our main result relies on a graphical representation of random choice rules. This directed graph is defined as follows. There is one node of this graph for each $A\in 2^{X}$. We index these nodes via $2^{X}$ and use the set $A$ to refer to the node indexed by $A$. There is a directed edge connecting $A$ to $B$ if $B=A\setminus\{x\}$ for some $x\in A$. Finally, we assign capacities to each edge. The edge capacity of the edge connecting $A$ and $A\setminus\{x\}$ is given by $q(x,A)$. We call this graphical representation the probability flow diagram. [^4] The probability flow diagram is originally due to Fiorini (2004).
$\emptyset$$\{x\}$$\{y\}$$\{z\}$$\{x,y\}$$\{x,z\}$$\{y,z\}$$\{x,y,z\}$$q(y,\{x,y,z\})$$q(x,\{x,y,z\})$$q(z,\{x,y,z\})$$q(y,\{x,y\})$$q(x,\{x,y\})$$q(z,\{x,z\})$$q(x,\{x,z\})$$q(z,\{y,z\})$$q(y,\{y,z\})$$q(x,\{x\})$$q(y,\{y\})$$q(z,\{z\})$
Figure 1: The probability flow diagram for the set $X_{n}=\{x,y,z\}$.
Definition 2.6.
A path $\rho$ is a finite sequence of sets $\{A_{i}\}_{i=0}^{n}$ such that $A_{i+1}\subsetneq A_{i}$ for all $i$, $A_{0}=X$, and $A_{|X|}=\emptyset$.
Observe that for every path, $|A_{i}\setminus A_{i+1}|=1$. In the probability flow diagram, each path from $X_{n}$ to $\emptyset$ bijectively corresponds to a preference. Each edge of the path corresponds to a unique input $(x,A)$ for the Möbius inverse. Using Theorem 2.1, a path that goes from $A_{0}$ to $A_{n}$ corresponds to the preference $A_{0}\setminus A_{1}\succ A_{1}\setminus A_{2}\succ\dots\succ A_{n-1}\setminus A_{n}$. To further describe this connection, if a random choice rule $p$ is induced by choice according to $\succ$, then each $q(x,A)$ on the path corresponding to $\succ$ is equal to one and each $q(y,B)$ not on the path corresponding to $\succ$ is equal to zero.
3 The Dimension of Preferences
In this section we present our main result and proof.
Theorem 3.1.
For any $n$, $\max_{M\in\mathcal{I}_{n}}|M|=(n-2)2^{n-1}+2$.
Before presenting our proof of Theorem 3.1, we first discuss some of its implications. Recall that for a given $X_{n}$, there are $n!$ strict preferences. We now compare the number of possible preferences with the maximal number of preferences in an identified model. Since preferences here are linear orders, the number of possible preferences is $n!$. According to Theorem 3.1, for reasonably large $n$, the size of a maximal identified model roughly doubles when adding a new alternative, while the size of the largest possible model, identified or otherwise, increases by a factor of $n$. In terms of ratios,
$\begin{split}\frac{(n-2)2^{n-1}+2}{n!}&=\frac{2^{n-1}}{n(n-1)(n-3)!}+\frac{2}{n!}\\ &\leq\frac{2^{n-1}}{(n-1)!}+\frac{2}{n!}\\ &\leq\frac{2^{n-1}}{(n-1)!}+\frac{2}{(n-1)!}\\ &=\frac{2^{n-1}+2}{(n-1)!}.\end{split}$ (3)
This tells us that the ratio of the maximal number of preferences in an identified model and the total number of preferences grows at a rate $\mathcal{O}(\frac{2^{n-1}}{(n-1)!})$. Clearly, as $n$ gets large, this ratio rapidly goes to zero. Thus, any identified model must rule out a large mass of preferences.
We now discuss the proof of Theorem 3.1. First, observe that there is an important connection between identified models and linear independence.
Definition 3.1.
A set of vectors $\{\mathbf{v}_{1},\dots,\mathbf{v}_{n}\}$ is linearly independent if any set of scalars $\{c_{1},\dots,c_{n}\}$ with $\sum_{i=1}^{n}c_{i}\mathbf{v}_{i}=0$ implies that $c_{i}=0$ for all $i$.
Let $p_{\succ}$ denote the real-valued vector, indexed by elements of the form $(x,A)$ with $x\in A$, which encodes the random choice rule induced by choice according to $\succ$. In other words, $p_{\succ}(x,A)=1$ if $x\succ y$ for all $y\in A\setminus\{x\}$ and $p_{\succ}(x,A)=0$ otherwise.
The following result is straightforward and known, but we reproduce it here for completeness.
Lemma 3.1.
A model $M$ is identified if and only if the set $\{p_{\succ}\}_{\succ\in M}$ is linearly independent.
Proof.
If $\{p_{\succ}\}_{\succ\in M}$ is not linearly independent, then for each $\succ\in M$, there is $c_{\succ}\in\mathbb{R}$ for which $\sum_{\succ\in M}c_{\succ}p_{\succ}=0$, where at least one $c_{\succ}\neq 0$. As each $p_{\succ}\geq 0$ componentwise, it follows that $\{\succ\in M:c_{\succ}>0\}\neq\varnothing$ and $\{\succ\in M:c_{\succ}<0\}\neq\varnothing$. Observe then that $\sum_{\{\succ\in M:c_{\succ}>0\}}c_{\succ}p_{\succ}=\sum_{\{\succ\in M:c_{\succ}<0\}}-c_{\succ}p_{\succ}$. Now, define $\nu\in\Delta(M)$ as $\nu(\succ^{\prime})=\frac{c_{\succ^{\prime}}}{\sum_{\{\succ\in M:c_{\succ}>0\}}c_{\succ}}$ when $c_{\succ^{\prime}}>0$, and $\nu(\succ^{\prime})=0$ otherwise. Similarly define $\nu^{\prime}\in\Delta(M)$ as $\nu(\succ^{\prime})=\frac{-c_{\succ^{\prime}}}{\sum_{\{\succ\in M:c_{\succ}>0\}}-c_{\succ}}$ when $c_{\succ^{\prime}}<0$ and $\nu(\succ^{\prime})=0$ otherwise. Observe that $\nu$ and $\nu^{\prime}$ have disjoint supports, but by construction $p_{\nu}=p_{\nu^{\prime}}$, so that $M$ is not identified. If $\{p_{\succ}\}_{\succ\in M}$ is linearly independent, it is easy to see that $M$ is identified: $p_{\nu}=p_{\nu^{\prime}}$ implies $\sum_{\succ\in M}(\nu(\succ)-\nu^{\prime}(\succ))p_{\succ}=0$, so that $\nu(\succ)-\nu^{\prime}(\succ)=0$ for all $\succ\in M$ by the definition of linear independence.∎
Now let $q_{\succ}$ denote the vector, indexed by elements of the form $(x,A)$ with $x\in A$, which encodes the Möbius inverse of the random choice rule induced by choice according to $\succ$. That is, $q_{\succ}(x,A)=1$ if $\succ\in L(x,A)$ and $q_{\succ}(x,A)=0$ otherwise. It is less obvious that a model $M$ is identified if and only if the set $\{q_{\succ}\}_{\succ\in M}$ is linearly independent. It essentially follows from Lemma 3.1 as the map carrying $p_{\succ}$ to $q_{\succ}$ is known to be linear and invertible. The following provides a formal proof.
Lemma 3.2.
A model $M$ is identified if and only if $\{q_{\succ}\}_{\succ\in M}$ is linearly independent.
Proof.
Rota (1964), Propositions 1 and 2, shows that there is a bijective and linear relationship between a Möbius inverse $q$ and its generating function $p$. This bijection is through Equation 2. Hence, there is an invertible linear map carrying generating functions $p$ to Möbius inverses $q$. The result follows as linear independence is preserved under invertible linear transformations. ∎
Our proof of Theorem 3.1 relies on the probability flow diagram with one small alteration. Specifically, we take the probability flow diagram and add an edge connecting $\emptyset$ to $X_{n}$ (and assign it an edge capacity equal to one). This motivates the following proposition.
Proposition 3.1.
Let $\{x^{i}\}_{i=1}^{I}$ be a collection of vectors for which for all $i,j$, $\sum_{k}x_{k}^{i}=\sum_{k}x_{k}^{j}\neq 0$. Then $\{x^{i}\}_{i=1}^{I}$ is linearly independent if and only if $\{(x^{i},1)\}_{i=1}^{I}$ is linearly independent, where $(x^{i},1)$ refers to the concatenation of the vector $x^{i}$ with a $1$.
Proof.
One direction is obvious (the linear independence of $\{x^{i}\}_{i=1}^{I}$ entails the linear independence of $\{(x^{i},1)\}_{i=1}^{I}$). For the other direction, suppose by means of contradiction that $\{(x^{i},1)\}_{i=1}^{K}$ is linearly independent but that there are $c^{i}$ not all 00 for which $\sum_{i}c^{i}x^{i}=0$. Let us suppose without loss that $\sum_{k}x^{i}_{k}=1$ for all $i$. Observe then that $\sum_{k}\sum_{i}c^{i}x^{i}_{k}=0$, and that $\sum_{k}\sum_{i}c^{i}x^{i}_{k}=\sum_{i}c^{i}\sum_{k}x^{i}_{k}=\sum_{i}c^{i}$. Consequently $\sum_{i}c^{i}=0$ and hence $\sum_{i}c^{i}(x^{i},1)=0$, contradicting linear independence of $\{(x^{i},1)\}_{i=1}^{I}$. ∎
In the proof of Theorem 3.1, we make use of two concepts from graph theory. First, a circuit is a sequence of directed edges where one edge’s starting node is the previous edge’s ending node with the last edge’s ending node being the same as the first edge’s starting node. In simpler terms, a circuit is a directed cycle. Second, the cyclomatic number of a directed graph corresponds to the minimal number of edges in the graph needed to be removed in order to remove all cycles from the graph. The cyclomatic number also corresponds to the circuit rank of the graph or the maximum number of independent circuits in the graph. We are now ready to proceed with the proof of Theorem 3.1.
Proof.
Terminology in this proof is as in Berge (2001). By Lemma 3.2 and Proposition 3.1, a model $M$ is identified if and only if the set of vectors $\{(q_{\succ},1)\}_{\succ\in M}$ is linearly independent. [^5] Observe that for each $\succ$, $\sum_{(x,A):x\in A}q_{\succ}(x,A)=n$, so that the condition in Lemma 3.2 is satisfied.
Now, consider the directed graph whose nodes are $2^{X_{n}}$ and has a directed edge from $A$ to $A\setminus\{x\}$ for any $x\in A$, and also has a directed edge from $\varnothing$ to $X_{n}$. We call this the appended probability flow diagram. The vector $(q_{\succ},1)$ then represents an indicator function of a subset of the appended probability flow diagram. The discussion following Definition 2.6 explains that $(q_{\succ},1)$ represents an indicator function of a circuit on the appended probability flow diagram, or an oriented loop (the $1$ reflects the edge going from $\varnothing$ to $X_{n}$).
By Theorem 3 of Chapter 4 of Berge (2001), the cyclomatic number of the appended probability flow diagram is equal to the dimension of the space spanned by indicator functions of circuits of the appended probability flow diagram. Now, the indicator function of any circuit of this diagram must be a linear combination of indicator functions of circuits corresponding to preferences. The reasoning is straightforward: by construction, every circuit must pass through the edge connecting $\varnothing$ to $X_{n}$. A circuit which passes through this edge only once clearly corresponds to a preference. A circuit passing through it $k$ times corresponds to a concatenation of $k$ circuits passing through this edge only once. Hence, the indicator function of the circuit passing through this edge $k$ times is the sum of the indicator functions of the $k$ circuits which pass through the edge only once, each of which correspond to a preference. Hence, the dimension of the circuit space of the appended probability flow diagram is the same as the dimension of the space spanned by $\{(q_{\succ},1)\}$ for all preferences $\succ$.
The cyclomatic number of a strongly connected graph, as our appended probability flow diagram is, according to Berge (2001), is defined by $E-N+1$, where $E$ is the number of edges and $N$ is the number of nodes. There are $2^{n}$ nodes in our appended probability flow diagram. There is one edge in our probability flow diagram for each $(x,A)$ with $x\in A$. This is given by $\sum_{i=1}^{n}i\binom{n}{i}$, which is equal to $n2^{n-1}$. [^6] The argument is standard, for each element of a set of size $n$, there are $2^{n-1}$ sets containing it. So $\sum_{i=1}^{n}i\binom{n}{i}$, which counts the sum of cardinalities of all subsets, must be given by summing the number $2^{n-1}$ across each element of the set of size $n$, resulting in $n2^{n-1}$. See Feller (1968), p.63. This means that our appended probability flow diagram has $n2^{n-1}+1$ edges, accounting for the edge going from $\varnothing$ to $X_{n}$. It then follows that the cyclomatic number of our appended probability flow diagram is given by the following:
$n2^{n-1}+1-2^{n}+1=(n-2)2^{n-1}+2.$ (4)
Thus, the space spanned by $(q_{\succ},1)$ has dimension $(n-2)2^{n-1}+2$. This tells us that for any model $M$, the dimension of the space spanned by $(q_{\succ},1)$ as $\succ\in M$ is at most $(n-2)2^{n-1}+2$, and that this bound is achieved for some model. In particular, any $M$ for which $(q_{\succ},1)$ is linearly independent must have $|M|\leq(n-2)2^{n-1}+2$ and this bound is achieved for some $M$. Again by Proposition 3.1, any $M$ for which $q_{\succ}$ is linearly independent must have $|M|\leq(n-2)2^{n-1}+2$ and this bound is achieved. Conclude by Lemma 3.2 that any identified model must have $|M|\leq(n-2)2^{n-1}+2$ and that this bound is achieved. ∎
The dimension of all random choice rules is the same as the dimension of the set of RUM random choice rules; see e.g. Corollary 6(ii) of Saito (2018) or Theorem 2 of Dogan and Yildiz (2022), where it is shown that the rational choice functions span the span of all stochastic choice functions. This immediately gives us the following corollary to Theorem 3.1, which is not so difficult to prove directly by simple combinatorial arguments.
Corollary 3.1.
The set of random choice rules on $X_{n}$ has dimension $(n-2)2^{n-1}+2$.
Corollary 3.1 obtains bounds on identified models even absent the RUM restriction.
4 A Sufficient Condition for Linear Independence
In this section, using the graphical intuition developed in the last section, we propose a new condition which guarantees that a model is identified. While this condition is sufficient for identification, we show by example that it fails to be necessary. We then compare this condition to two other uniqueness conditions in the random utility literature. We now introduce our new condition.
Definition 4.1.
We say that a model $M$ is edge decomposable if for every nonempty submodel $N\subseteq M$ there exists a preference $\succ\in N$ and a tuple $(x,A)$ with $x\in A$ such that $N\cap L(x,A)=\{\succ\}$.
Our argument often leverage a form of edge decomposability which takes the form of an induction argument:
Definition 4.2.
We say that a model $M$ is sequentially edge decomposable if it can be enumerated as $\{\succ_{1},\ldots,\succ_{k}\}$ so that for each $i\in\{1,\ldots,k\}$, there is a tuple $(x,A)$ with $x\in A$ such that $\{\succ_{i},\ldots,\succ_{k}\}\cap L(x,A)=\{\succ_{i}\}$.
A simple induction argument guarantees the following, which we will often use without mention.
Proposition 4.1.
A model $M$ satisfies edge decomposabilitiy iff it satisfies sequential edge decomposability.
Keeping in mind Theorem 2.1, when a model is edge decomposable, we are able to do the following to recover the unique distribution over preferences in our model which induces the data. First, as $M$ is a (weak) submodel of $M$, there exists some preference $\succ$ and pair $(x,A)$ such that {}$\{\succ\}=M\cap L(x,A)$. By Theorem 2.1, this tells us that $q(x,A)=\nu(L(x,A))=\nu(\succ)$. Now consider the submodel which is given by $N=M\setminus\{\succ\}$. By edge decomposability, we know that there is some $\succ^{\prime}\in N$ such that {}$\{\succ^{\prime}\}=N\cap L(y,B)$ for some $y\in B$. It then follows that $\nu(\succ^{\prime})=q(y,B)$ if $\succ\not\in L(y,B)$ and $\nu(\succ^{\prime})=q(y,B)-\nu(\succ)$ if $\succ\in L(y,B)$. Since $M$ is finite, we can repeated this argument to get full recovery of our distribution $\nu$ over preferences in $M$, and so every edge decomposable model $M$ is identified. The following proposition formally restates this observation.
Proposition 4.2.
If a model $M$ is edge decomposable, then M is identified.
We now interpret our definition of edge decomposability in terms of the probability flow diagram. Recall that every pair $(x,A)$ with $x\in A$ is associated with a specific edge of the probability flow diagram. Specifically, $q(x,A)$ is assigned as the edge capacity of the edge associated with $(x,A)$. Edge decomposability then says, given a model $M$, for each submodel $N$ there exists some path associated with $\succ\in N$ such that this path has some edge which is unique to it among paths associated with preferences $\succ^{\prime}\in N$. Following a similar process to what was described in the last paragraph, this means that when a graph is written as a combination of flows along paths associated with preferences in $\succ$ in a edge decomposable model $M$, we can always uniquely decompose this graph into flow weights on paths associated with preferences in $M$ [^7] Further, we can think of the analogue of edge decomposability for graphs other than the probability flow diagram. In this case, edge decomposability still guarantees that we are able to uniquely decompose our graph into a series of flow into our edge decomposable set of paths. While edge decomposability may be sufficient for identification of a model $M$, it fails to be necessary. We show this through Example 4.1 and Figure 2.
Example 4.1.
Consider the environment given by $X_{n}=\{a,b,c,d,e,f,g,h\}$. We now provide a model with eight preferences which is not edge decomposable and yet is identified. These preferences are represented on the probability flow diagram of Figure 2. The paths and preferences are as follows.
  1. 1.
    Black path - $\succ_{1}$ $f\succ g\succ d\succ h\succ c\succ e\succ a\succ b$
  2. 2.
    Red path - $\succ_{2}$ $h\succ g\succ e\succ f\succ b\succ d\succ a\succ c$
  3. 3.
    Pink path - $\succ_{3}$ $f\succ g\succ h\succ e\succ d\succ c\succ b\succ a$
  4. 4.
    Blue path - $\succ_{4}$ $h\succ g\succ f\succ d\succ c\succ e\succ b\succ a$
  5. 5.
    Aqua path - $\succ_{5}$ $g\succ f\succ d\succ h\succ e\succ b\succ a\succ c$
  6. 6.
    Green path - $\succ_{6}$ $g\succ h\succ f\succ d\succ e\succ b\succ c\succ a$
  7. 7.
    Purple path - $\succ_{7}$ $g\succ f\succ h\succ e\succ b\succ d\succ c\succ a$
  8. 8.
    Orange path - $\succ_{8}$ $g\succ h\succ e\succ f\succ d\succ c\succ b\succ a$
We now proceed to show that $M=\{\succ_{1},\succ_{2},\succ_{3},\succ_{4},\succ_{5},\succ_{6},\succ_{7},\succ_{8}\}$ is an identified model despite not being edge decomposable. It follows immediately from Figure 2 that $M$ is not edge decomposable as each edge covered by $M$ is covered by two paths. Now we show identification by directly calculating the values of $\nu$. To begin, we start by identifying the probability weight on the preferences $\succ_{2}$, $\succ_{4}$, and $\succ_{8}$.
$\begin{split}\nu(\succ_{2})+\nu(\succ_{4})&=q(h,\{a,b,c,d,e,f,g,h\})\\ \nu(\succ_{2})+\nu(\succ_{8})&=q(e,\{a,b,c,d,e,f\})\\ \nu(\succ_{4})+\nu(\succ_{8})&=q(b,\{a,b\})\end{split}$ (5)
This gives us the following.
$\begin{split}\nu(\succ_{2})&=\frac{q(h,\{a,b,c,d,e,f,g,h\})+q(e,\{a,b,c,d,e,f\})-q(b,\{a,b\})}{2}\\ \nu(\succ_{4})&=\frac{q(h,\{a,b,c,d,e,f,g,h\})-q(e,\{a,b,c,d,e,f\})+q(b,\{a,b\})}{2}\\ \nu(\succ_{8})&=\frac{-q(h,\{a,b,c,d,e,f,g,h\})+q(e,\{a,b,c,d,e,f\})+q(b,\{a,b\})}{2}\end{split}$ (6)
The probability weight on each other preference is then given by the following.
$\begin{split}\nu(\succ_{1})&=q(c,\{a,b,c,e\})-\nu(\succ_{4})\\ \nu(\succ_{6})&=q(f,\{a,b,c,d,e,f\})-\nu(\succ_{4})\\ \nu(\succ_{3})&=q(a,\{a,b\})-\nu(\succ_{1})\\ \nu(\succ_{5})&=q(e,\{a,b,c,e\})-\nu(\succ_{6})\\ \nu(\succ_{7})&=q(c,\{a,c\})-\nu(\succ_{6})\\ \end{split}$ (7)
Note that each line of Equation 7 is written only using variables that have been identified from Equation 6 or higher lines in Equation 7. Equations 6 and 7 tell us we can recover our distribution over preferences $\nu$ directly from the data for model $M$, and so $M$ is identified despite not being edge decomposable.
$\emptyset$$\{c\}$$\{a\}$$\{b\}$$\{a,c\}$$\{a,b\}$$\{a,c,d\}$$\{a,b,c\}$$\{a,b,e\}$$\{a,b,c,d\}$$\{a,b,c,e\}$$\{a,b,c,d,f\}$$\{a,b,c,d,e\}$$\{a,b,c,e,h\}$$\{a,b,c,d,e,f\}$$\{a,b,c,d,e,h\}$$\{a,b,c,d,e,f,g\}$$\{a,b,c,d,e,f,h\}$$\{a,b,c,d,e,g,h\}$$\{a,b,c,d,e,f,g,h\}$
Figure 2: The graphical representation of Example 4.1. Each differently colored path corresponds to a different preference in our counterexample.
As Example 4.1 demonstrates, edge decomposability is not necessary for identification. However, edge decomposability is a condition that is relatively easy to use. By that we mean that it is easy to constructively create edge decomposable models. Simply start with a preference $\succ$ and consider the model $N=\{\succ\}$. Given $N$, find a pair $(x,A)$ with $x\in A$ such that $N\cap L(x,A)=0$. Choose any $\succ^{\prime}\in L(x,A)$ and create a larger model $N\cup\{\succ^{\prime}\}$. We can repeat this process until no such pairs $(x,A)$ exist. Once no such pairs exist, we have a maximal edge decomposable model.
4.1 Relationship to Other Uniqueness Conditions
We now compare edge decomposability to two other uniqueness conditions in the random utility literature. Our focus is on the uniqueness conditions introduced in Turansick (2022) and Apesteguia et al. (2017). While both of these restrictions recover uniqueness in different ways, they both work by placing restrictions on the preferences in the induced support of their model. Our main goal is to compare these restrictions on supports to our edge decomposability condition on models. We will show that the supports induced by both Turansick (2022) and Apesteguia et al. (2017) are subsets of the possible supports/models induced by edge decomposability.
To begin, we review the setup and characterization of Turansick (2022). The goal of Turansick (2022) is to characterize when the model $M=\mathcal{L}(X_{n})$ admits a unique representation. Note that this is different from our definition of identification. In the proof of Theorem 1 from Turansick (2022), it is shown that a random choice rule $p$ has a unique representation if and only if every supported path in the probability flow diagram has some edge unique to that path among supported paths. Here, by supported paths, we mean paths where every edge on that path has a strictly positive edge capacity. Restating this in term of preferences, this means that $p$ has a unique support if and only if one of the following conditions hold for each preference.
  1. 1.
    $\succ\in L(x,A)$ and $q(x,A)=0$
  2. 2.
    $\exists(x,A)$ with $x\in A$ such that ${}L(x,A)\cap\{\succ^{\prime}|\succ^{\prime}\in L(y,B)\implies q(y,B)>0\}=\{\succ\}$
Relating this back to our definition of edge decomposability, let $M=\{\succ^{\prime}|\succ^{\prime}\in L(y,B)\implies q(y,B)>0\}$. Then it is the case that the result of Turansick (2022) is asking that for every $\succ\in M$, there exists some $(x,A)$ with $x\in A$ such that $M\cap L(x,A)=\{\succ\}$. The condition of Turansick (2022) is obviously weakly stronger than edge decomposability. We now show through an example, related to the counterexample of Fishburn (1998), that it is strictly stronger than edge decomposability.
Example 4.2.
Consider the model $M$ given by the following three preferences.
  1. 1.
    $a\succ b\succ c\succ d=\succ_{1}$
  2. 2.
    $b\succ a\succ d\succ c=\succ_{2}$
  3. 3.
    $a\succ b\succ d\succ c=\succ_{3}$
Observe the following.
  1. 1.
    $M\cap L(a,\{a,b,c,d\})=\{\succ_{1},\succ_{3}\}$
  2. 2.
    $M\cap L(b,\{b,c,d\})=\{\succ_{1},\succ_{3}\}$
  3. 3.
    $M\cap L(c,\{c,d\})=\{\succ_{2},\succ_{3}\}$
  4. 4.
    $M\cap L(d,\{d\})=\{\succ_{2},\succ_{3}\}$
This means that $\succ_{3}$ never has a set $L(x,A)$ with $\succ_{3}\in L(x,A)$ that is unique to $\succ_{3}$ among preferences in $M$. However, as $M\cap L(b,\{a,b,c,d\})=\succ_{2}$ and $M\cap L(d,\{c,d\})=\{\succ_{1}\}$, it is easy to see that $M$ is edge decomposable.
[…]
Definition 4.3.
We say that a distribution over preferences $\nu$ satisfies the single-crossing property if the support of $\nu$ can be ordered in such a way that for all $x\rhd y$, $x\succ_{i}y$ implies $x\succ_{j}y$ for all $j\geq i$.
Unlike previously, it is not immediately obvious that the supports of SCRUM representations are edge decomposable. In order to see that they are, note the following.
Proposition 4.3.
[…]
Proof.
If $\nu$ has a single preference in its support, we are done. Suppose otherwise that $n\geq 2$, so that that $\succ_{1}$ and $\succ_{2}$ are in the support of $\nu$. Since $\succ_{1}$ and $\succ_{2}$ differ, this means that there exists some pair $(x,y)$ such that $x\succ_{1}y$ and $y\succ_{2}x$. Since $\nu$ satisfies the single crossing property, for all $i\geq 2$, we have that $y\succ_{i}x$. Let $A$ denote the set such that $\succ_{1}\in L(x,A)$. Since $x\succ_{1}y$, $y\in A$ and thus $L(x,A)\cap\{\succ_{1},\dots,\succ_{n}\}=\{\succ_{1}\}$. The result now continues by induction. ∎
It follows from Proposition 4.3 that the supports of SCRUM representations satisfy edge decomposability. Specifically, we know that every subset of a SCRUM support is itself the support of some SCRUM representation. It then follows that the lowest ranked preference in the support is the unique element of some $L(x,A)$ among preferences in the support. In other words, once the support of a SCRUM representation is pinned down, we can recursively find the probability weights on the preferences by looking at the lowest and then the next lowest ranked preferences in the support. We now show through an example that SCRUM supports fail to capture every edge decomposable model.
Example 4.3.
[…]
  1. 1.
    $a\succ b\succ c\succ d\succ f\succ e=\succ_{1}$
  2. 2.
    $a\succ b\succ d\succ c\succ e\succ f=\succ_{2}$
  3. 3.
    $b\succ a\succ c\succ d\succ e\succ f=\succ_{3}$
We now proceed with some case work.
  1. 1.
    Suppose $a\rhd b$. Then it is the case that, abusing notation, if $M$ can satisfy the single-crossing property, $\succ_{1}\rhd\succ_{3}$ and $\succ_{2}\rhd\succ 3$.
    1. (a)
      Suppose $c\rhd d$. Then it is the case that $\succ_{1}\rhd\succ_{2}$ and $\succ_{3}\rhd\succ_{2}$. However, given that $a\rhd b$ and $\succ_{2}\rhd\succ_{3}$, $c\rhd d$ cannot be the case.
    2. (b)
      Suppose $d\rhd c$. This means that $\succ_{2}\rhd\succ_{1}$ which further means that $e\rhd f$. However, as $a\rhd b$ and $\succ_{1}\rhd\succ_{3}$, we also have that $f\rhd e$, and so $d\rhd c$ cannot be the case.
  2. 2.
    Suppose $b\rhd a$. This means that $\succ_{3}\rhd\succ_{1}$ and $\succ_{3}\rhd\succ_{2}$. This gives us that $a\rhd b$, $c\rhd d$, and $e\rhd f$. As $c\succ_{1}d$, $d\succ_{2}c$, $f\succ_{1}e$, and $e\succ_{2}f$, we have that $\succ_{1}\rhd\succ_{2}\rhd\succ_{1}$, which is a cycle. Thus $b\rhd a$ cannot be the case.
The above case work tells us that it can neither be the case that $a\rhd b$ nor $b\rhd a$, and thus $\{\succ_{1},\succ_{2},\succ_{3}\}$ cannot be a SCRUM support. However, as $\succ_{1}$ is the only preference to rank $f\succ e$, $\succ_{2}$ is the only preference to rank $d\succ c$, and $\succ_{3}$ is the only preference to rank $b\succ a$, the model $M=\{\succ_{1},\succ_{2},\succ_{3}\}$ is edge decomposable.
[…]
5 The Maximal Identified Models for $n=4$
We now show how our Theorem 3.1 can be used to verify that certain sets of preferences are maximal sets of linearly independent preferences. Note that all such sets then act as a basis for the set of random choice rules. This follows for the same reason as discussed prior to Corollary 3.1. Our goal now is to characterize every maximal set of linearly independent preferences when $n=4$. Using our Theorem 3.1, this tells us that we are looking for sets with $18$ preferences of the total possible $24$. In order to do this, we first return to the example of Fishburn (1998) which shows that random utility is not identified.
Example 5.1 (Fishburn (1998)).
Let $X=\{a,b,c,d\}$. Consider the following probability distributions over linear orders on $X$.
$\nu_{1}(\succ)=\begin{cases}\frac{1}{2}&\text{if }\succ\in\{a\succ b\succ c\succ d,b\succ a\succ d\succ c\}\\ 0&\text{otherwise}\end{cases}$
$\nu_{2}(\succ)=\begin{cases}\frac{1}{2}&\text{if }\succ\in\{a\succ b\succ d\succ c,b\succ a\succ c\succ d\}\\ 0&\text{otherwise}\end{cases}$
These two probability distributions induce the same random choice rule.
It turns out that the two distributions over preferences in Fishburn’s example, $\nu_{1}$ and $\nu_{2}$, are the (only) extreme points of the set of distributions which can induce the random choice rule implied by Example 5.1. This means that if our model $M$ only contains any three of the four preferences in Example 5.1, then we regain identification. To see this, suppose that our model is $M=\{a\succ b\succ c\succ d,b\succ a\succ d\succ c,a\succ b\succ d\succ c\}$. This model $M$ is edge decomposable since $M\cap L(c,\{c,d\})=\{a\succ b\succ c\succ d\}$, $M\cap L(a,\{a,c,d\})=\{b\succ a\succ d\succ c\}$, and removal of $a\succ b\succ c\succ d$ leaves $a\succ b\succ d\succ c$ as the only preference in $M\cap L(a,\{a,b,c,d\})$ and removal of $b\succ a\succ d\succ c$ leaves $a\succ b\succ d\succ c$ as the only preference in $M\cap L(d,\{c,d\})$. This means that our model $M$ is identified.
We can permute the names of alternatives in Example 5.1 and come up with another set of four preferences which lacks identification. Observe that the four preferences in the example are $\{a\succ b\succ c\succ d,a\succ b\succ d\succ c,b\succ a\succ c\succ d,b\succ a\succ d\succ c\}$. That is, these are precisely the preference which rank each member of $\{a,b\}$ above each member of $\{c,d\}$. We can write this set of preferences as $\{a,b\}\succ\{c,d\}$.
Consequently, there are multiple analogues of Fishburn’s example with four alternatives, indexed by the lower set. As in Example 5.1, one way is with $\{a,b\}\succ\{c,d\}$. But another would be $\{a,c\}\succ\{b,d\}$, and still another would be $\{b,c\}\succ\{a,d\}$. All of these sets of preferences are disjoint. There are exactly six total ways of choosing two alternatives to be ranked above the other two alternatives. These six different rankings can be indexed by the lower set. In the case of Example 5.1, this lower set is $\{c,d\}$. Now observe that the set of preferences used in Example 5.1 is exactly $L(c,\{c,d\})\cup L(d,\{c,d\})$. We just argued that we can recover identification for a model $M\subseteq L(c,\{c,d\})\cup L(d,\{c,d\})$ if $|M|=3$.
Recall that, by Theorem 3.1, we are looking for sets with $18$ preferences. There are six different binary sets when $n=4$ and thus six different analogues of Fishburn’s example. Further, in each one of these variations on Example 5.1, we can recover identification by removing one of the four preferences, leaving us with three preferences for each variant of Example 5.1. This gives us a total of $18$ preferences, the maximal number of linearly independent preferences when $n=4$. This condition turns out to be characteristic of every maximal set of linearly independent preferences when $n=4$.
Proposition 5.1.
Suppose $n=4$. Then a model $M$ is a maximal set of linearly independent preferences if and only if for each set $A$ with $|A|=2$ and each $x,y\in A$ such that $x\neq y$, we have that $|M\cap(L(x,A)\cup L(y,A))|=3$.
Proof.
We begin by considering a model $M$ satisfying the conditions of Proposition 5.1. Let $\varnothing\neq M^{\prime}\subseteq M$. Then there exists some $\{c,d\}$ such that $M^{\prime}\cap\{a,b\}\succ\{c,d\}\neq\varnothing$. By hypothesis, $|M^{\prime}\cap\{a,b\}\succ\{c,d\}|\leq 3$. The argument following Example 5.1 demonstrates that (by permutation) that for any $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$ of cardinality at most three, there are $(x,A)$ with $x\in A$ and ′′$\succ^{\prime\prime}\in M^{\prime\prime}$ for which $M^{\prime\prime}\cap L(x,A)=\{\succ^{\prime\prime}\}$. In particular, this holds for some ′′$\succ^{\prime\prime}\in M^{\prime}\cap\{a,b\}\succ\{c,d\}$. Without loss of generality, suppose that $a\succ b\succ c\succ d$ is not a member of $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$. There are three cases. If $a\succ b\succ c\succ d$ is in $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$, then we can choose our tuple to be $(b,\{b,c,d\})$. If $b\succ a\succ c\succ d$ is in $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$, then we can choose $(c,\{c,d\})$ to be our tuple. If neither $a\succ b\succ c\succ d$ nor $b\succ a\succ c\succ d$ is in $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$, then $M^{\prime\prime}\subseteq\{a,b\}\succ\{c,d\}$ is exactly $\{b\succ a\succ d\succ c\}$ and we can choose either $(b,\{b,c,d\})$ or $(d,\{d,c\})$ as our tuple. For any of these sets, and any $\{x,y\}\neq\{c,d\}$, $\{x,y\}\succ\{z,w\}\cap L(x,A)=\varnothing$. Consequently $M^{\prime}\cap L(x,A)=\{\succ^{\prime\prime}\}$.
Conversely, let $M$ be a maximal set of linearly independent preferences. By Theorem 3.1, $|M|=18$. Suppose to a contradiction that there is some $A=\{x,y\}$, where $x\neq y$, for which $|M\cap(L(x,A)\cup L(y,A))|\neq 3$. If $|M\cap(L(x,A)\cup L(y,A))|=4$, then the four preferences in $M\cap(L(x,A)\cup L(y,A))$ are exactly the four preferences in Example 5.1 (up to permutation) and thus $M$ is not an identified model. If instead $|M\cap(L(x,A)\cup L(y,A))|\leq 2$, then since there are six binary sets and $18$ preferences in $M$, by the pigeonhole principle there must be some set $B=\{c,d\}$ with $|B|=2$ such that $|M\cap(L(c,B)\cup L(d,B))|=4$ for $x,y\in B$ and $x\neq y$. The four preferences in $M\cap(L(c,B)\cup L(d,B))$ are exactly the four preferences in Example  5.1 (up to permutation) and thus $M$ is not an identified model. ∎
Proposition 5.1 exactly characterizes every preference basis for random choice rules when $n=4$. Unfortunately, the logic used in the proof of Proposition 5.1 does not extend obviously to $n>4$. Proposition 5.1 also tells us exactly how many maximal identified models there are when $n=4$. Since there are six different binary sets, and we are choosing one of four preferences associated with that binary set to not be in our model, there are $4404^{6}=61,$ 440 different maximal identified models when $n=4$. Compare this to exactly one maximal identified model when $n=3$ (every model is identified in that case).
6 Related Literature
Our paper is primarily related to two strands of literature. The first strand is the one which studies uniqueness and identification in random utility models. Study of the random utility model goes back to Block and Marschak (1959) and Falmagne (1978). Barberá and Pattanaik (1986) and Fishburn (1998) note that the random utility model is in general not identified when there are at least four alternatives available. Much of the literature studying identification aims to recover uniqueness by refining the random utility model. The random expected utility model of Gul and Pesendorfer (2006) recovers uniqueness by restricting to expected utility functions while studying choice over lotteries. As discussed earlier, the single-crossing random utility model of Apesteguia et al. (2017) is able to recover uniqueness by asking that the support of their representation satisfy the single-crossing property with respect to some exogenous order. Yildiz (2023) extends this line of thought and characterizes exactly which models of choice can be identified through a similar progressivity condition. Another approach to identification is through the use of stronger data. Lu (2019) is able to recover both beliefs and preferences when stochastic choice data and information sources are observed. Dardanoni et al. (2020) is able to recover preferences as well as cognitive heterogeneity with data that connects agents’ choices across menus. An alternative approach is taken by Turansick (2022). Turansick (2022) simply asks which realizations of standard random choice rules admit a unique random utility representation. Azrieli and Rehbeck (2022) studies stochastic choice and random utility when we are unable to condition choice on the menu of available alternatives. They find that in this setting that the random utility model is also unidentified. Kashaev et al. (2024) studies separable stochastic choice and finds that separability can be characterized by each agent having well defined marginal choices if and only if the set of feasible choice functions (i.e. preferences) are linearly independent.
The second strand of literature our paper is related to is the strand which uses graph theoretic tools to study the random utility model and more generally stochastic choice. To our knowledge, this strand of literature was started by Fiorini (2004) who used the observation that preferences can be represented as flows on the probability flow diagram in order to provide an alternative proof of the characterization of Falmagne (1978). Turansick (2022) studies conditions on the probability flow diagram which characterize when random choice rules have a unique random utility representation. Chang et al. (2022) studies which preferences have adjacent paths in the probability flow diagram in order to say when random-coefficient models are good approximations of the random utility model. Chambers et al. (Forthcoming) extends the probability flow diagram to multiple dimension in order to study choice across multiple people or time periods. Finally, Kono et al. (2023) use graph theoretic tools to study the random utility model when choice probabilities of some alternatives are unobserved.